OpenStudy (anonymous):

Find the exact value sin11pi/2

OpenStudy (anonymous):

Split it up into \[ \sin(8\pi/2 + 3\pi/2) = \sin(4\pi + 3\pi/2) \]Then use: \[ \sin(x+y) = \sin(x)\cos(y)+\sin(y)\cos(x) \]

OpenStudy (anonymous):

Why do you split it up?

OpenStudy (anonymous):

Because you can find the exact value for angles \(4\pi\) and \(3\pi/2\) very easily.

OpenStudy (campbell_st):

because of the cyclic nature of sin(x) you just need to find \[\sin(\frac{3\pi}{2})\]